One posotive number is 5 times another number. The difference between the two numbers is 1132, find the numbers?

Thank you

Answer by macston(5194)

Let x=smaller number
5x-x=1132
4x=1132
x=283
5x=5(283)=1415 ANSWER: The numbers are 283 and 1415.
CHECK:
1415-283=1132
1132=1132